A hot water tank has 10cm of foam (k= 0.02 Wm^-1 K^-1) insulation surrounding the 60°C tank which is placed in 20°C room. The tank has surface area of 3.5 m^2 . Find the thermal energy lost over a one day period. This will equal the standby loss of tank. Ignore air films.


1
Expert's answer
2021-01-31T03:43:01-0500

Rate of heat flow is given by "\frac{dQ}{dt} = KA\Delta T"

where K is thermal conductivity, A is the area, dt is the time, dQ is the heat and "\Delta T" is the temperature difference.

Here we have, "K = 0.02, A = 3.5, \Delta T = 60-20=40, t = 1day = 24*60*60=86400s"


Putting all values, Amount of heat flow is given by

"dQ = KA\Delta T (dt) = 0.02 \times 3.5\times40\times86400 = 241920 J = 241.92 KJ"